# Discovering overlapping communities in multi-layer directed networks

### v1

@@ 4.1 Oracle case
The form of $U_{r}=\Pi_{r}U_{r}(\mathcal{I}_{r},:)$ is known as ideal simplex, where the rows of $U_{r}$ locate in a simplex in $\mathbb{R}^{K}$ with vertices being the $K$ rows of $U_{r}(\mathcal{I}_{r},:)$. Such ideal simplex has been observed in single layer mixed membership community detection [25, 17, 19, 18, 26], topic modeling [27, 28], and latent class analysis [29, 30]. Since $U_{r}(\mathcal{I}_{r},:)$ is a $K\times K$ full rank matrix, we have $\Pi_{r}=U_{r}U^{-1}_{r}(\mathcal{I}_{r},:)$, which indicates that we can exactly recover the row membership matrix through $U_{r}U^{-1}_{r}(\mathcal{I}_{r},:)$ as long as the index set $\mathcal{I}_{r}$ is available. Benefited from the simplex structure $U_{r}=\Pi_{r}U_{r}(\mathcal{I}_{r},:)$, applying the Successive Projection Algorithm (SPA) [20, 14, 21] to all rows of $U_{r}$ with $K$ clusters can exactly hunt for the $K$ vertices (i.e., the $K$ rows of $U_{r}(\mathcal{I}_{r},:)$) in the simplex. Similarly, $U_{c}=\Pi_{c}U_{c}(\mathcal{I}_{c},:)$ also forms an ideal simplex structure and applying the SPA to $U_{c}$ can exactly recover $\mathcal{I}_{c}$. The following four-stage algorithm called ideal Co-clustering by Sequential Projection on the Debiased Sum of Squared matrices (ideal CSPDSoS) summarizes the above analysis. Input: $\{\Omega_{l}\}^{L}_{l=1}$ and $K$. Output: $\Pi_{r}$ and $\Pi_{c}$.
@@ 5 Main results
Assumption 1 says that the sparsity parameter $\rho$ should diminish at a rate slower than $\sqrt{\frac{\mathrm{log}(n+L)}{n^{2}L}}$, aligning with the sparsity condition specified in Theorem 1 of [7]. We also assume that the smallest singular values of $\sum_{l\in[L]}B_{l}B^{\prime}_{l}$ and $\sum_{l\in[L]}B^{\prime}_{l}B_{l}$ satisfy a linear growth respective to the layers $L$.
@@ 6.1 Synthetic data
Algorithms for comparison. While [1] proposed an efficient method for community detection in multi-layer ScBM for non-overlapping multi-layer directed networks, it is unable to estimate nodes’ mixed memberships in overlapping multi-layer directed networks. Therefore, similar to [7, 1], we consider the following two approaches for comparison:
@@
Evaluation metrics. Hamming Error and Relative Error serve as metrics for evaluating the performance of various methods. Hamming Error is calculated as $\mathrm{max}\left(\frac{\mathrm{min}_{\mathcal{P}\in S}\|\hat{\Pi}_{r}\mathcal{P}-\Pi_{r}\|_{1}}{n},\frac{\mathrm{min}_{\mathcal{P}\in S}\|\hat{\Pi}_{c}\mathcal{P}-\Pi_{c}\|_{1}}{n}\right)$, where $S$ represents the set of all $K\times K$ permutation matrices. This metric ranges from 0 to 1, with a lower value indicating superior performance. Relative Error, on the other hand, is defined as $\mathrm{max}\left(\frac{\mathrm{min}_{\mathcal{P}\in S}\|\hat{\Pi}_{r}\mathcal{P}-\Pi_{r}\|_{F}}{n},\frac{\mathrm{min}_{\mathcal{P}\in S}\|\hat{\Pi}_{c}\mathcal{P}-\Pi_{c}\|_{F}}{n}\right)$ and is a non-negative value. The closer it is to zero, the better the performance.
@@
(c) Experiment 2
(d) Experiment 2
(e) Experiment 3
(f) Experiment 3
@@
Results. By Figure 2, we see that our proposed CSPDSoS approach exhibits superior performance in estimating $\Pi_{r}$ and $\Pi_{c}$ when we increase $n$, $L$, or the overall sparsity $\rho$. This observation validates our theoretical findings. Additionally, it is noteworthy that CSPDSoS consistently outperforms CSPSoS, and both methods significantly surpass CSPSum, indicating the effectiveness of our proposed approach.
@@ 6.2 Real data
We apply our CSPDSoS to the FAO Multiplex Trade Network (FAOMTN), a real-world multi-layer directed network accessible for download at https://manliodedomenico.com/data.php in this subsection. This dataset constitutes a comprehensive global network of food imports and exports, featuring multiple layers that represent different products. The nodes in this network denote individual countries, while the edges within each layer represent the import/export connections between countries for a particular product. The dataset, compiled in 2010, encompasses 364 products and 213 countries [31], where we have consolidated China and China-mainland into a single category labeled “China” to ensure consistency. We have arranged the layers in decreasing order according to their total edge count and selected the 30 most densely connected layers (i.e., $n=213,L=30$). For a comprehensive listing of the 30 products utilized in this study, please refer to Table 1.
Initially, the data was weighted, but we have simplified it to an unweighted format whenever a trade relationship exists between countries. To enhance interpretability, we have set $K=6$, representing the number of continents excluding Antarctica. For a country $i$, if $\mathrm{max}_{k\in[K]}\hat{\Pi}_{r}(i,k)\leq 0.5$, it is classified as a highly mixed export country, suggesting a diversified economy with exports to various countries. Similarly, if $\mathrm{max}_{k\in[K]}\hat{\Pi}_{c}(i,k)\leq 0.5$, it is deemed a highly mixed import country. Figure 3 shows the lists of these highly mixed export (import) countries and their estimated mixed memberships obtained from CSPDSoS. The term $\mathrm{argmax}_{k\in[K]}\hat{\Pi}_{r}(i,k)$ (and $\mathrm{argmax}_{k\in[K]}\hat{\Pi}_{c}(i,k)$) refers to the home base row (and column) community of node $i$ for $i\in[n]$. Figure 4 presents the home base community assignments returned by the CSPDSoS method for three selected layers, offering a clear visualization of the identified communities. Figure 5 presents a visual representation of the estimated home base row and column communities for the 213 countries considered in this study. Notably, the structure of the row (export) communities exhibits distinct differences compared to the column (import) communities, indicating a variance in trade patterns across exporting and importing nations.

### v2

@@ 4.1 Oracle case
The form of $U_{r}=\Pi_{r}U_{r}(\mathcal{I}_{r},:)$ is known as ideal simplex, where the rows of $U_{r}$ locate in a simplex in $\mathbb{R}^{K}$ with vertices being the $K$ rows of $U_{r}(\mathcal{I}_{r},:)$. Such ideal simplex has been observed in single layer mixed membership community detection [58, 8, 52, 50, 57], topic modeling [59, 60], and latent class analysis [61, 62]. Since $U_{r}(\mathcal{I}_{r},:)$ is a $K\times K$ full rank matrix, we have $\Pi_{r}=U_{r}U^{-1}_{r}(\mathcal{I}_{r},:)$, which indicates that we can exactly recover the row membership matrix through $U_{r}U^{-1}_{r}(\mathcal{I}_{r},:)$ as long as the index set $\mathcal{I}_{r}$ is available. Benefited from the simplex structure $U_{r}=\Pi_{r}U_{r}(\mathcal{I}_{r},:)$, applying the Successive Projection Algorithm (SPA) [53, 45, 54] to all rows of $U_{r}$ with $K$ clusters can exactly hunt for the $K$ vertices (i.e., the $K$ rows of $U_{r}(\mathcal{I}_{r},:)$) in the simplex. The details of SPA are provided in Algorithm 1. Upon execution on $U_{r}$, SPA accepts $U_{r}$ and $K$ as inputs (with $M=U_{r}$ and $r=K$ in Algorithm 1). SPA iteratively chooses $K$ rows from the eigenvector matrix $U_{r}$ that possess the highest Frobenius norms. Each selected row prompts SPA to eliminate its contribution within the matrix, thereby ensuring the orthogonality of subsequent selections to previously chosen rows. Consequently, a collection of $K$ pure nodes $\mathcal{I}$, is obtained. For deeper insights into SPA, please refer to [8, 50, 62].
Algorithm 1 SPA
1:Matrix $M\in\mathbb{R}^{n\times K}$ and integer $r\leq n$.
2:Set of indices $\mathcal{I}\subseteq[n]$.
3:Initialize: ${}\mathcal{I}=\{\}$.
4:For $k\in[r]$ do:
5:       $k_{*}=\mathrm{argmax}_{i\in[n]}\|M(i,:)\|_{F}$.
6:      $M\leftarrow M-M\frac{M^{\prime}(k_{*},:)M(k_{*},:)}{\|M(k_{*},:)\|^{2}_{F}}$.
7:      =$\mathcal{I}=\mathcal{I}\cup\{k_{*}\}$.
8:end for
Similarly, $U_{c}=\Pi_{c}U_{c}(\mathcal{I}_{c},:)$ also forms an ideal simplex structure and applying the SPA to $U_{c}$ can exactly recover $\mathcal{I}_{c}$. The following four-stage algorithm called ideal Co-clustering by Sequential Projection on the Debiased Sum of Squared matrices (ideal CSPDSoS) summarizes the above analysis. Input: $\{\Omega_{l}\}^{L}_{l=1}$ and $K$. Output: $\Pi_{r}$ and $\Pi_{c}$.
@@ 4.2 Real case
Remark 2.
Our CSPDSoS method, primarily designed to estimate asymmetric mixed memberships of nodes in multi-layer directed networks, can be effortlessly adapted for application in multi-layer bipartite networks. For example, consider a multi-layer bipartite network comprising $L$ layers, where $A^{bi}_{l}\in\{0,1\}^{n_{r}\times n_{c}}$ for $l\in[L]$, with $n_{r}$ and $n_{c}$ representing the number of row and column nodes, respectively [63, 57]. Here, $A^{bi}_{l}$ denotes the bipartite adjacency matrix for the $l$-th layer. By merely substituting $A_{l}$ in Algorithm 2 with $A^{bi}_{l}$, we can estimate mixed memberships for both row and column nodes in multi-layer bipartite networks.
@@ 5 Main results
Assumption 1 says that the sparsity parameter $\rho$ should diminish at a rate slower than $\sqrt{\frac{\mathrm{log}(n+L)}{n^{2}L}}$, aligning with the sparsity condition specified in Theorem 1 of [28]. Similar to Assumption 1.(b) in [28] and Assumption 2 in [1], we also assume that the smallest singular values of $\sum_{l\in[L]}B_{l}B^{\prime}_{l}$ and $\sum_{l\in[L]}B^{\prime}_{l}B_{l}$ satisfy a linear growth respective to the layers $L$.
@@ 6.1 Synthetic data
Algorithms for comparison. While [1] proposed an efficient method for community detection in multi-layer ScBM for non-overlapping multi-layer directed networks, it is unable to estimate nodes’ mixed memberships in overlapping multi-layer directed networks. Meanwhile, as evidenced by Table 1, except for our CSPDSoS algorithm, as far as we know, there is no method specifically designed for estimating mixed memberships in overlapping multi-layer directed networks. Therefore, similar to [28, 1, 57], we consider the following approaches for comparison:
@@
  • 3.
    GeoNMF, SVM-cone-DCMMSB (SVM-cD for short), and Mixed-SCORE: These three algorithms, initially presented in [64], [65], and [50] respectively, are tailored for estimating nodes’ mixed memberships in single-layer undirected networks. To adapt these algorithms for detecting overlapping communities in multi-layer directed networks, certain modifications are necessary. Inspired by our CSPDSoS algorithm and the DSoG algorithm in [1], we use the aggregation matrices $S_{r}$ and $S_{c}$ to substitute for their adjacency matrix to estimate the row membership matrix $\Pi_{r}$ and the column membership matrix $\Pi_{c}$, respectively.
Remark 3.
We should emphasize that the MASE [22], SoS-Debias [28], CAMSBM [29], and DSoG [1] algorithms are tailored for detecting nodes’ non-overlapping communities in multi-layer networks but can’t be adapted for estimating mixed memberships in such networks. Thus, we exclude them and their variants from our comparison. To our knowledge, the three approaches CSPDSoS, CSPSum, and CSPSoS developed in this paper are the only algorithms specifically designed for estimating mixed memberships in multi-layer directed networks. Given that GeoNMF, SVM-cD, and Mixed-SCORE are initially designed for estimating nodes’ mixed memberships in single-layer undirected networks, we’ve adapted them for multi-layer directed networks to provide a more comprehensive comparison with our CSPDSoS method.
Evaluation Metrics. When the true membership matrices $\Pi_{r}$ and $\Pi_{c}$ are known, we consider the following three metrics:
  • 1.
    Hamming error: This metric is defined as $\mathrm{max}\left(\frac{\mathrm{min}_{\mathcal{P}\in\mathcal{G}}\|\hat{\Pi}_{r}\mathcal{P}-\Pi_{r}\|_{1}}{n},\frac{\mathrm{min}_{\mathcal{P}\in\mathcal{G}}\|\hat{\Pi}_{c}\mathcal{P}-\Pi_{c}\|_{1}}{n}\right)$, where $\mathcal{G}$ represents the set of all $K\times K$ permutation matrices. Consistent with the measure used in Theorem 1, which establishes the theoretical upper bound for per-node error rates through the $l_{1}$ difference between true and estimated mixed memberships for both row and column clusters, the Hamming error ranges from 0 to 1, with lower values indicating better performance.
  • 2.
    Relative error: This metric is defined as $\mathrm{max}\left(\frac{\mathrm{min}_{\mathcal{P}\in\mathcal{G}}\|\hat{\Pi}_{r}\mathcal{P}-\Pi_{r}\|_{F}}{\|\Pi_{r}\|_{F}},\frac{\mathrm{min}_{\mathcal{P}\in\mathcal{G}}\|\hat{\Pi}_{c}\mathcal{P}-\Pi_{c}\|_{F}}{\|\Pi_{c}\|_{F}}\right)$. Unlike the Hamming error, which measures the $l_{1}$ difference, the Relative error assesses the relative $l_{2}$ difference between the true and estimated membership matrices for both row and column communities. It is worth noting that both Hamming and Relative errors have been used in [65, 8, 52, 57, 50, 61] to evaluate the discrepancy between the true and estimated mixed membership matrices. A deeper comprehension of Hamming error and Relative error can be found in [57].
  • 3.
    Overlapping Normalized Mutual Information (ONMI): Introduced by [66], this metric extends NMI [67, 68] to evaluate overlapping community detection. It ranges from 0 to 1, with higher values indicating better performance. Detailed computations of ONMI can be found in [66, 7, 69], which we omit here. It is important to note that ONMI applies only to binary membership matrices [65, 7]. However, the membership matrices $\Pi_{r},\Pi_{c},\hat{\Pi}_{r}$, and $\hat{\Pi}_{c}$ considered in this study are continuous. Therefore, to utilize ONMI, we convert these matrices into their binary counterparts. For $\Pi_{r}$, we define its binary version as $\Pi^{0}_{r}$, where $\Pi^{0}_{r}(i,k)=1$ if $\Pi_{r}(i,k)\geq\delta$ and 0 otherwise, for $i\in[n]$ and $k\in[K]$. Here, $\delta$ is a threshold value set to $\frac{1}{K}$ in this study. Similarly, we define $\Pi^{0}_{c},\hat{\Pi}^{0}_{r}$, and $\hat{\Pi}^{0}_{c}$ as the binary versions of $\Pi_{c},\hat{\Pi}_{r}$, and $\hat{\Pi}_{c}$, respectively. Let $\mathrm{ONMI}(\Pi^{0}_{r},\hat{\Pi}^{0}_{r})$ represent the ONMI between $\Pi^{0}_{r}$ and $\hat{\Pi}^{0}_{r}$, and $\mathrm{ONMI}(\Pi^{0}_{c},\hat{\Pi}^{0}_{c})$ represent the ONMI between $\Pi^{0}_{c}$ and $\hat{\Pi}^{0}_{c}$. To simultaneously assess the true and estimated binary membership matrices for both row and column communities, we define the ONMI used in this paper as the minimum of $\mathrm{ONMI}(\Pi^{0}_{r},\hat{\Pi}^{0}_{r})$ and $\mathrm{ONMI}(\Pi^{0}_{c},\hat{\Pi}^{0}_{c})$, given that a higher ONMI value indicates better performance.
@@
(c) Experiment 1
(d) Experiment 1
(e) Experiment 2
(f) Experiment 2
(g) Experiment 2
(h) Experiment 2
(i) Experiment 3
(j) Experiment 3
(k) Experiment 3
(l) Experiment 3
@@
Results. The experimental results of Experiments 1-3 are displayed in Fig. 3 and they provide a rigorous validation of the proposed CSPDSoS algorithm for overlapping community detection in multi-layer directed networks, while also highlighting the performance of the CSPSoS, GeoNMF, SVM-cD, and Mixed-SCORE algorithms adapted for this task. Experiment 1 demonstrates that as the number of nodes $n$ increases, the CSPDSoS, CSPSoS, GeoNMF, SVM-cD, and Mixed-SCORE algorithms significantly outperform CSPSum in terms of Hamming error, Relative error, and ONMI, indicating more accurate estimation of the row and column membership matrices. This trend aligns with the theoretical prediction that increasing the number of nodes improves detection accuracy. Experiment 2 shows that increasing the number of layers $L$ in the multi-layer directed network further enhances the performance of CSPDSoS compared to the other algorithms, underscoring the value of additional layers in providing more information for community detection. Experiment 3 reveals that as the overall sparsity parameter $\rho$ decreases, representing denser networks, all algorithms achieve better performance, with lower Hamming and Relative errors, and higher ONMI. Particularly, our CSPDSoS enjoy competitive performances with CSPSoS, GeoNMF, SVM-cD, and Mixed-SCORE. This indicates that the proposed CSPDSoS algorithm is particularly effective in handling networks with denser connections. By panel (d) of this figure, we see that all methods are capable of detecting communities in a multi-layer directed network comprising 5000 nodes and 10 layers within 40 seconds. Overall, the consistent improvements across all three experiments not only confirm the theoretical analysis but also demonstrate the superior performance of the CSPDSoS, CSPSoS, SVM-cD, GeoNMF, and Mixed-SCORE methods for overlapping community detection in multi-layer directed networks.
Remark 4.
Though GeoNMF, SVM-cD, and Mixed-SCORE have competitive performances with our CSPDSoS algorithm, it is important to note that we have modified the original GeoNMF, SVM-cD, and Mixed-SCORE algorithms from single-layer undirected networks to multi-layer directed networks, as done in Remark 3. Therefore, the versions of GeoNMF, SVM-cD, and Mixed-SCORE used in this article are new methods tailored for multi-layer directed networks. Additionally, though these modified methods can detect mixed memberships for multi-layer directed networks generated from our multi-layer MM-ScBM model, we cannot obtain their theoretical analysis immediately. Instead, building their theoretical guarantees would require new models or techniques, which we leave as future work.
@@ 6.2 Real data
We apply our CSPDSoS to the FAO Multiplex Trade Network (FAOMTN), a real-world multi-layer directed network accessible for download at https://manliodedomenico.com/data.php in this subsection. This dataset constitutes a comprehensive global network of food imports and exports, featuring multiple layers that represent different products. The nodes in this network denote individual countries, while the edges within each layer represent the import/export connections between countries for a particular product. The dataset, compiled in 2010, encompasses 364 products and 213 countries [70], where we have consolidated China and China-mainland into a single category labeled “China” to ensure consistency. We have arranged the layers in decreasing order according to their total import/export value of the product and selected the top 30 products with the highest trading volume (i.e., $n=213,L=30$). For a comprehensive listing of the 30 products utilized in this study, please refer to Table 2. Meanwhile, Tables 3 and 4 present the 25 countries with the smallest and largest import/export trade volumes, respectively, ranked based on their trade volume size among 213 countries. Initially, the data was weighted, but we have simplified it to an unweighted format whenever the trading value of the product between countries reaches or exceeds 100. After preprocessing the data, for the FAOMTN multi-layer directed network considered in this article, we have $A_{l}\in\{0,1\}^{213\times 213}$ for $l\in[30]$.
Table 3: The 25 countries with the smallest import/export trade volumes. The numbers in parentheses represent the ranking of the trade volumes, where 1 indicates the smallest (i.e., ranked last among 213 countries) and 25 indicates the 25th smallest (i.e., ranked 189th among 213 countries).
Mayotte (1) Andorra (2) Falkland Islands Malvinas (3) Comoros (4) Saint Pierre and Miquelon (5)
Tuvalu (6) Palau (7) Kiribati (8) Montserrat (9) Nauru (10)
Cook Islands (11) Eritrea (12) Marshall Islands (13) Micronesia, Federated States of (14) Niue (15)
Saint Helena (16) Somalia (17) Guam (18) Samoa (19) Brunei Darussalam (20)
Cayman Islands (21) Grenada (22) American Samoa (23) British Virgin Islands (24) Equatorial Guinea (25)
Table 4: The 25 countries with the largest import/export trade volumes. The numbers in parentheses represent the ranking of the trade volumes, where 1 indicates the largest (i.e., ranked first among 213 countries) and 25 indicates the 25th largest (i.e., ranked 25th among 213 countries).
United States (1) Germany (2) China (3) Netherlands (4) France (5)
Brazil (6) United Kingdom (7) Italy (8) Japan (9) Indonesia (10)
Canada (11) Belgium (12) Spain (13) Malaysia (14) Argentina (15)
Mexico (16) Russia (17) Australia (18) Thailand (19) Denmark (20)
Poland (21) India (22) Korea, Republic of (23) Switzerland (24) Austria (25)
Since the actual row and column memberships of the data are unknown, it is not feasible to assess the efficacy of the six methods employed on synthetic data. Instead, we consider the estimated row and column membership matrices generated by our CSPDSoS algorithm as the ground truth. By computing the Hamming error, Relative error, and ONMI for the remaining five methods as the number of communities increases, we aim to investigate how closely these methods compare to our CSPDSoS algorithm in analyzing this real-world dataset. Fig. 4 displays the results. We observe that as the number of communities $K$ increases, the estimated mixed memberships from the other five methods diverge more significantly from the estimated mixed membership matrices obtained from our CSPDSoS algorithm.
Figure 4: Comparison of different methods on the FAOMTN dataset using estimated row and column membership matrices from our CSPDSoS algorithm as ground truth. Plots show the Hamming error, Relative error, and ONMI for the remaining five methods as the number of communities increases.
Figure 5: Overall asymmetric degree measured by Hamming error, Relative error, and ONMI between $\hat{\Pi}_{r}$ and $\hat{\Pi}_{c}$ returned by our CSPDSoS algorithm when $K$ increases for the FAOMTN dataset.
Since the true mixed membership matrices for the FAOMTN datasets remain unknown, we rely solely on our CSPDSoS algorithm for subsequent analysis. Given that FAOMTN is a multi-layer directed network, we are particularly interested in determining whether there is a significant asymmetry in the row and column community memberships. To answer this question, we compare the discrepancies between $\hat{\Pi}_{r}$ and $\hat{\Pi}_{c}$ generated by our CSPDSoS algorithm as the number of communities $K$ increases. This allows us to investigate the overall asymmetry between row and column communities. Fig. 5 presents the results. It is evident that there is a notable asymmetry between the row and column mixed communities in the FAOMTN network, as indicated by the large Hamming and Relative errors, coupled with the relatively small ONMI values observed between $\hat{\Pi}_{r}$ and $\hat{\Pi}_{c}$. Recall that FAOMTN is a worldwide product import/export network, results in Fig. 5 imply significant differences in the community structures between global import and export trades. From an economic network perspective, this asymmetry suggests that the clustering of countries into communities differs notably depending on whether one is looking at imports or exports.
Figure 6: Asymmetry score against the number of communities $K$ for the FAOMTN dataset. The red and blue dashed lines on the y-axis indicate the thresholds of $y=0.5$ and $y=0.0001$, respectively. Nodes with Asymmetry scores exceeding 0.5 are denoted by red solid circles, those with Asymmetry scores greater than 0.0001 but less than or equal to 0.5 are represented by black hollow circles, and those with Asymmetry scores less than or equal to 0.0001 are depicted by blue solid circles.
However, measuring the difference between $\hat{\Pi}_{r}$ and $\hat{\Pi}_{c}$ can only evaluate the overall asymmetric between row and column mixed communities. To find which node has the largest asymmetric pattern, we adopt the index Asymmetry score introduced in [9]. For $i\in[n]$, this index is defined as below:
$a_{i}(K)=\|\hat{U}_{r}(i,:)-\hat{U}_{c}(i,:)\|_{F},$
where $\hat{U}_{r}$ and $\hat{U}_{c}$ are eigenvector matrix of $S_{r}$ and $S_{c}$, respectively. If the network is undirected, the Asymmetry score for any node is 0. A node with a high Asymmetry score exhibits distinct sending and receiving patterns [9]. Based on the results presented in Fig. 6, we have the following observations:
  • 1.
    Only a few countries have Asymmetry scores no larger than 0.0001, indicating no significant difference between their import and export trades. In particular, we find 14 countries with an Asymmetry score no larger than 0.0001. These countries include Comoros, Nauru, Eritrea, Marshall Islands, Falkland Islands Malvinas, Cook Islands, Montserrat, Saint Pierre and Miquelon, Micronesia, Federated States of, Palau, Kiribati, Tuvalu, Mayotte, Andorra. This suggests that these countries’ import and export activities are relatively balanced, possibly indicating economic self-sufficiency or a match between the demand for their imported and exported goods. Alternatively, it could indicate that these countries have minimal or almost no import/export relationships with other countries because we observe that all these 14 countries are listed in Table 3, i.e., their trade volumes are among the smallest in the 213 countries considered.
  • 2.
    For the majority of countries, the Asymmetry scores fall within the range of $(0.0001,0.5]$, suggesting a moderate difference between their import and export trades. This implies that while there is some disparity between what these countries import and export, the difference is not overwhelming. This moderate asymmetry may reflect a diversified economy where both imports and exports play a significant role.
  • 3.
    Only a few countries have Asymmetry scores that surpass 0.5, indicating a pronounced asymmetry between their import and export trades. Specifically, when $K=2$, it is exclusive to Germany; when $K$ is 3 or 4, this characteristic is observed in France, Germany, and the Netherlands; when $K=5$, it applies to Belgium, France, Germany, Italy, Netherlands, Poland, United Kingdom, and United States; when $K=6$, it is seen in Belgium, France, Germany, Italy, Netherlands, United Kingdom, and United States; when $K=7$, it includes Belgium, France, Germany, Italy, Netherlands, Poland, Spain, United Kingdom, and United States; when $K=8$, it pertains to Belgium, Canada, France, Germany, Italy, Netherlands, Spain, United Kingdom, and United States; and when $K=9$, it encompasses Belgium, Canada, China, France, Germany, Italy, Netherlands, Poland, Spain, United Kingdom, and United States. These countries are often major players in the global economy, with substantial import and export volumes, and they may experience significant trade surpluses or deficits. This observation aligns with the data presented in Table 4, as their trade volumes rank among the highest among all countries. It can be concluded that countries with small import/export volumes typically have very low Asymmetry scores, indicating no significant difference between their imports and exports, while countries with large import/export volumes often have high Asymmetry scores, indicating a significant difference between their imports and exports.
Table 5: $\eta$ and $\sigma^{2}$ for all estimated communities for the FAOMTN network.
$C_{r,1}$ $C_{r,2}$ $C_{r,3}$ $C_{r,4}$ $C_{r,5}$ $C_{r,6}$ $C_{c,1}$ $C_{c,2}$ $C_{c,3}$ $C_{c,4}$ $C_{c,5}$ $C_{c,6}$
$\eta$ 0.0719 0.0826 0.0598 0.0717 0.0832 0.6309 0.0364 0.0558 0.0833 0.1759 0.2473 0.4013
$\sigma^{2}$ 0.0245 0.0223 0.0105 0.0131 0.0157 0.1691 0.0060 0.0119 0.0234 0.0632 0.0757 0.1285
To enhance interpretability, we have set $K=6$, representing the number of continents considered for trade analysis, excluding Antarctica for this particular study. We then apply our CSPDSoS approach to the FAOMTN dataset, assuming 6 row/column (import/export) communities. This results in the estimated mixed membership matrices $\hat{\Pi}_{r}\in[0,1]^{213\times 6}$ for import trade and $\hat{\Pi}_{c}\in[0,1]^{213\times 6}$ for export trade. Subsequently, we delve into a deeper analysis of these two estimated mixed membership matrices. Recall that the $n\times 1$ membership vector $\hat{\Pi}_{r}(:,k)$ represents the extent to which each country belongs to the $k$-th estimated row (import) community $C_{r,k}$ for $k\in[K]$. We are interested in examining the likelihood of countries belonging to this community, as well as assessing the diversity within this community. We denote the mean of this vector by $\eta$ and its variance by $\sigma^{2}$. Specifically, we compute $\eta$ to measure the likelihood of countries belonging to this community, whereas $\sigma^{2}$ serves as an indicator of the diversity in these mixed memberships. Based on the results recorded in Table 5, we have the following observations:
  • 1.
    Row (import) communities: The $\eta$ values range from 0.0598 to 0.6309. This indicates that some communities (e.g., $C_{r,3}$ with $\eta$=0.0598) have relatively low mean membership values, suggesting that these 213 countries are less likely to belong exclusively to these communities. In contrast, communities like $C_{r,6}$ ($\eta$=0.6309) have higher mean membership values, indicating that these countries are more likely to belong to this community. Meanwhile, the $\sigma^{2}$ values range from 0.0105 to 0.1691. Lower variance values (e.g.,$C_{r,3}$ with $\sigma^{2}=0.0105$) indicate that countries within the community have more homogeneous membership values, while higher variance values (e.g., $C_{r,6}$ with $\sigma^{2}=0.1691$) suggest higher membership diversity.
  • 2.
    Column (export) communities: The $\eta$ values range from 0.0364 to 0.4013. Similar to the import communities, some export communities have lower mean membership values (e.g., $C_{c,1}$ with $\eta$=0.0364), while others have higher values (e.g., $C_{c,6}$ with $\eta$=0.4013). Meanwhile, the $\sigma^{2}$ values range from 0.0060 to 0.1285. Similar to import communities, some export communities exhibit lower variance (e.g., $C_{c,1}$ with $\sigma^{2}=0.0060$), indicating homogeneity, while others show higher variance (e.g., $C_{c,6}$ with $\sigma^{2}=0.1285$), indicating diversity.
  • 3.
    Comparison between import and export Communities: The mean membership values $\eta$ for import communities are generally higher than those for export communities, suggesting that countries tend to have stronger affiliations with import communities. The variance in membership values $\sigma^{2}$ for import communities is also generally higher than for export communities, indicating greater diversity in membership affiliations within import communities. These findings suggest that import communities have stronger and more diverse affiliations among countries.
For country $i$, if $\mathrm{max}_{k\in[K]}\hat{\Pi}_{r}(i,k)\leq 0.4$, it is classified as a highly mixed import country, suggesting a diversified economy with imports from various countries. Similarly, if $\mathrm{max}_{k\in[K]}\hat{\Pi}_{c}(i,k)\leq 0.4$, it is deemed a highly mixed export country. Fig. 7 shows the lists of these highly mixed import (export) countries and their estimated mixed memberships obtained from CSPDSoS. The results indicate that these countries exhibit low membership values across all import/export communities, suggesting a lack of strong preference for any specific one. This indicates a diversified economy, where they import and export various goods from multiple sources and regions. This pattern may arise from their small size or economic diversification strategies, which aligns with the observation that many of these countries are small and not heavily dependent on a particular good or region. Instead, they trade with multiple partners to satisfy their economic needs, and some may have implemented diversification strategies to minimize dependency on a single export or import. Fig. 8 displays the mixed memberships for countries with the smallest trade volumes, as listed in Table 3. Like Fig. 7, the heatmap reveals that these countries have low membership values in all import and export communities.
Fig. 9 displays the mixed memberships for the top 25 countries with the largest import/export trade volumes in the FAOMTN dataset. The heatmaps reveal interesting patterns of membership distributions across different import (row) and export (column) communities. These patterns provide insights into the complex trade dependencies and integration of countries within the global economy. In detail, we have the following observations:
  • 1.
    In the import communities, countries like the United States, Germany, Netherlands, France, Italy, India, China, Brazil, Japan, Indonesia, Malaysia, Argentina, and Mexico show strong membership values predominantly in one import community, suggesting these countries have dominant trade relationships with a particular economic bloc. In particular, the pure nodes are found to be United States, Germany, Netherlands, France, Italy, and India. On the other hand, countries like the United Kingdom, Canada, Belgium, Spain, Russia, Denmark, Poland, Switzerland, and Austria exhibit more balanced memberships across multiple import communities, indicating their trade relations are diversified and spread across various economic groups.
  • 2.
    Similarly, in the export communities, certain countries like Germany, Netherlands, France, Brazil, United Kingdom, Argentina, Mexico, Thailand, and India display concentrated memberships in one export community, reflecting their dominance in a specific export market. In particular, the pure nodes are found to be Germany, Netherlands, France, and United Kingdom. Conversely, countries such as the United States, China, Italy, Japan, Indonesia, Canada, Belgium, Spain, Malaysia, Russia, Australia, Denmark, Poland, Korea, Republic of, Switzerland, and Austria show more diverse distributed memberships across multiple export communities, suggesting they are active players in multiple export markets.
Let the term $\mathrm{argmax}_{k\in[K]}\hat{\Pi}_{r}(i,k)$ (and $\mathrm{argmax}_{k\in[K]}\hat{\Pi}_{c}(i,k)$) refer to the home base row (and column) community of node $i$ for $i\in[n]$. Fig. 10 presents the home base community assignments returned by the CSPDSoS method for three selected layers, offering a clear visualization of the identified communities. We observe that (a) in the “Soybeans” layer and the “Crude material” layer, there are many highly mixed countries in the import communities, indicated by the prevalence of green squares, while there are only a few highly mixed countries in the export communities; (b) for the “Food prep nes” layer, there are several highly mixed countries in both import and export communities; (c) across all three layers, the import communities differ significantly from the export communities, reflecting the asymmetric nature of trade relationships. Thus, this figure illustrates the complex and often asymmetric community structures in the trade network.
Fig. 11 and Fig. 12 present global maps of estimated mixed membership intensity in the six import communities and the six export communities for all 213 countries, respectively. These maps offer valuable insights into the dominant trade positions of countries within each import and export community. For example, in the import community $C_{r,1}$ (where the mixed memberships of countries belonging to this community are recorded in $\hat{\Pi}_{r}(:,1)$, the first column of $\hat{\Pi}_{r}$), it is evident that Germany stands out with the darkest color, signifying its dominant position in this community. This observation aligns with the results shown in Fig. 9, where Germany is identified as a pure node in $C_{r,1}$, further reinforcing its central role in this import community. The colors in the last panel of Fig. 11 (representing $C_{r,6}$) are generally darker than those in the other five panels, indicating that the membership values of countries belonging to $C_{r,6}$ are typically larger than those in the other import communities. This finding is consistent with the statistical summary in Table 5, which shows that the mean membership value ($\eta$) for $C_{r,6}$ is the highest among all import communities. Moreover, the maps reveal that countries with closer geographical locations tend to have similar membership intensities within each import community. For instance, European (or Asian) countries exhibit closely aligned memberships in all import communities, suggesting that these regions share similar import relationships and dependencies. A similar pattern emerges in the global map of export communities (Fig. 12). Germany, Netherlands, France, and United Kingdom stand out as dominant players in various export communities. As with the import communities, the membership intensities of countries within export communities also tend to cluster geographically.